[Halld-jlab] Hall D Staff Meeting Notes, October 6,. 2014
Mark Ito
marki at jlab.org
Mon Oct 6 21:21:05 EDT 2014
Folks,
Please find the notes below and at
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_6,_2014
-- Mark
_______________________________
Private:Hall D Staff Meeting, October 6, 2014
Present: Fernando Barbosa, Eugene Chudakov (chair), Mark Dalton,
Alexandre Deur, Hovanes Egiyan, Sergey Furletov, Mark Ito, David
Lawrence, Yi Qiang, Elton Smith, Sascha Somov, Simon Taylor, Beni Zihlmann
Eugene
* Appraisals will be initiated today. They are due by Oct. 14.
o With uncertainty due to the upcoming continuing resolution
Physics had decided no promotions for now
* 12 GeV budget: project is done
o all construction accounts close
o one exception is account for pre-ops
o Eugene will issue guidance on how to use it
o should be used for paying non-Hall-D groups at the lab
* Next Long-Range Plan meeting in April
* New PAC members: Frank Maas, Dean [?], Renee Fatemi, Werner Boeglin
* Post-doc search: choice sent to HR
David
* work needed on translation tables
* auto-transfer to tape library needs directory handling logic included
* online monitoring: getting help from offline folks: Paul Mattione,
Sean Dobbs, Kei Moriya
* need to modify read-out list to get DAQ parameters into data stream
* EPICS variables to data stream needs work
Sergey F.
* DAQ tested with CDC, FDC, BCAL, TOF, ST, FCAL in read-out
* read-out lists are scattered, need to consolidate
* highest rate so far 100 Hz with test pulser
o Sascha reports DAQ crashes on DMA from F125s
* monitoring of rates, occupancy being developed
* DAQ has gone at 17 kHz with no VME transfers
* Eugene asked about tagger hall readout
o Sascha: hodoscope has been read-out, worked with Nathan Sparks
o S: microscope, jumper settings for addresses for FADCs wrong,
Sergey P. is fixing
* DAQ settings and pedestals can be set for all F250s and F125s,
settings dumped to file
Sascha
* working on global trigger for BCAL
o using LED to fire trigger
o CDC had been added
o failed at 400 Hz
* finalizing read-out list for all detectors
* want set-ups to be able to run with triggers and crates
* next plan: use same global energy for FCAL
o will first check in local mode
o timing issues in combining BCAL and FCAL
* next area of work: pair spectrometer
o trigger through GTP, course counters only
* Eugene: can specific crates be excluded from read-out based in
trigger type?
o S: in principle yes, not implemented
* Vladimir Berdnikov has seen PS signals, looked at SiPM noise for
each bias group
o read-out is done
Fernando
* the Hall of Shame is expanding
o rack moved on upstream platform
* cables are all settled except for long-term network switch
o will need a PC to control microscope electronics
* Eugene: we plan to remove the tagger system on Friday
Yi
* solenoid
o power supply
+ 2 spare controller boards came in
+ 1 repaired
+ makes 4 total
o slew rate slower than that set
+ suspicion: either connect of board or fault in daughter board
+ will order a spare daughter board
+ can do a test at low current
+ should not impede operations
o Eugene: helium in magnet for a month now
o Jonathan Creel adjusted parameters on JT valves, smoother
operation now
o Eugene: flow controller for leads on the way
* active collimator
o high-speed EPICS readout implemented
o analyzer can read from ROOT file
Mark I.
* set-up and ran farm jobs for simulation of commissioning
o in support of work by Simon, Paul, Sean
* Eugene: Gagik Gavalian has developed an EVIO-to-ROOT converter for
Hall B
o we will follow-up to see if we can use it
* start counter
o need TDC readout
+ TDC read-out not a priority
+ yet TDC is the principle read-out method for this detector
o need self-trigger
+ start counter trigger not a priority
+ yet start counter trigger useful for cosmic testing of many
other detector subsystem e. g., CDC, BCAL (radial track path
enhancement), FDC (in coincidence with TOF to give
horizontal cosmics) and may be useful for beam data as well
Simon
* continuing work on commissioning geometry
* CDC alignment
o parameter cut-off necessary in fits
o more data with BCAL as trigger needed
+ last large run did not have CDC times, neither from TDCs nor
from FADCs
Mark D.
* working on BCAL data analysis, low-level reconstruction code
o Eugene: what is z-resolution?
+ with LED data, using ADC timing, about 3 cm
Beni
* F125 programming
o will average over 16 samples to determine pedestal
o one ROC has the new firmware
* why flashes cause DAQ to crash?
o sometimes one ROC does not see trigger
o need to maintain sync in those cases
o DMA buffer read-out error only problem on ROC-FDC-10
* working on CDC monitoring histograms
Alexandre
* implement the run plan
o procedures need work
* OSLs being placed in tagger hall[?]
Elton
* implementing BCAL config for LED pulsing through trigger supervisor,
getting help from Beni
* cell-phone-carrying personnel cross-training on Wed.
* safety walk-through Thu. at 9:30 am
* Eugene has a cell phone charging station, Yi will take it off his
hands and install it in the proper location
Hovanes
* beam profiler is back in position after survey
o it is downstream of FCAL, and connected, no counts yet
* halo counters installed, need cabling to electronics
* working on overview gui for EPICS, summary of important EPICS
channel all in one place
* installing some PDU for devices that may need remote power cycling
* moving test set-up for amorphous radiator harp from EEL to tagger
hall in case motion is needed for the run
* getting access to Accelerator EPICS GUIs
* Fernando asked about keys for cabinets, Tom had a plan for keybox,
Hovanes will check on status
--
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://mailman.jlab.org/pipermail/halld-jlab/attachments/20141006/84256bb9/attachment.html
More information about the Halld-jlab
mailing list